A chemist needs 500 milliliters of a 20% acid and 80% water mix for a chemistry experiment. The chemist combines X milliliters of a 10% acid and 90% water mix and Y milliliters of a 30% acid and 70% water mix to make the 20% acid and 80% water mix.
A. write a linear system that represents the situation. MY ANSWER: x+y= 500 -My teacher did part a in class.
B. How many milliliters of the 10% acid and 90% water mix and the 30% acid and 70% water mix are combined to make the 20% acid and the 80% water mix?
C. The chemists also needs 500 milliliters of a 15% acid and 85% water mix. Does the chemist need more of the 10% acid and 90% water mix than the 30% acid and 70% water mix to make this new mix? Explain.

Answer by stanbon(75887)

A chemist needs 500 milliliters of a 20% acid and 80% water mix for a chemistry experiment.
The chemist combines X milliliters of a 10% acid and 90% water mix and Y milliliters of a 30% acid and 70% water mix to make the 20% acid and 80% water mix.
------

A. write a linear system that represents the situation. MY ANSWER: x+y= 500 -My teacher did part a in class.
Note: 10% fluid + 30% fluid = 100 milliliters
acid + acid = acid
------------------------------------
Quantity:: n + r = 100 ml
Acid:::: 10n+30r = 20*100
-----
n + r = 100
n + 3r = 2*100
--
Subtract and solve for "r":: 2r = 100 ; r = 50ml (amt. of 30% solution needed)
Then n = 100 - 50 = 50ml (amt. of 10% solution needed)
--------------
-----

B. How many milliliters of the 10% acid and 90% water mix and the 30% acid and 70% water mix are combined to make the 20% acid and the 80% water mix?
Worked above.
----------------

C. The chemists also needs 500 milliliters of a 15% acid mix. Does the chemist need more of the 10% acid mix than the 30% acid mix to make this new mix? Explain.
Quantity: n + r = 500
Acid:::: 10n+30r = 15*500
--------------------
n + r = 500
n + 3r = 1.5*500
---
2r = 0.5*500
r = 0.5*250 = 125ml (amt. of 30% needed)
n = 500-125 = 375ml (amt. of 10% needed)
==================
Cheers,
Stan H.
------------